OpenStudy (anonymous):

Solve the system of equations. -2x + 3y = 8 3x – y = -5

OpenStudy (anonymous):

multiply second equation by 3 add the two together

OpenStudy (anonymous):

Try this: Multiply the second by 3. Add it the the first one. tell us what do you get.

OpenStudy (anonymous):

how do i set that up?

OpenStudy (anonymous):

Like this: 1st :-2x+3y=8 2nd*(3) : 9x-3y=-15

OpenStudy (anonymous):

Now add them. Tell us what do you get.

OpenStudy (anonymous):

\[-2x + 3y = 8\]\[ 3x – y = -5\] \[-2x+3y=8\]\[3(3x-y)=3(-5)\] \[-2x+3y=8\]\[9x-3y=-15\] add to get \[7x=-7\]solve for \(x\)\[7x=-7\implies x=-1\]

OpenStudy (theviper):

y=3x+5 Substitute this value in eq 1

OpenStudy (anonymous):

thankss